Adjusted Trial Balance Cash $ 3,500 8,000 3,600 75,000 Prepaid insurance $12,000 600 30,000 14,100 Common stock. 7,500 102,200 4,200 Miscellaneous income.. Salaries expense. Rent expense. Insurance expense. Depreciation expense.. . 42,800 12,900 1,800 8,000 8,800 8,800 $171,900 $171,900

Prepare a detailed Balance Sheet and also calculate the following ratios:

A) Current Ratio

B) Working Capital

C) Assets Turnover

D) Fixed Assets Turnover

E) A/R Turnover

F) Debt Ratio
Answer #1
Income Statement
Particulars Amount($) Amount($)
Revenues
Service fee Earned 102200
Miscellaneous Income 4200
Total Revenues 106400
Less: Expenses
Salaries Expenses 42800
Rent Expenses 12900
Insurance Expenses 1800
Depreciation Expenses 8000
Income Tax Expenses 8800
Total Expenses 74300
Net Income 32100
Retained Earnings Statement
Particulars Amount($) Amount($)
Retained Earnings Jan 1, 20xx 14100
Net income 32100
Less: Dividends 7500
Retained Earnings 24600
Retained earnings Dec31, 20xx 38700
Stockholders Equity
Particulars Amount($)
Common stock 30000
Retained Earnings 38700
Total Stockholders Equity 68700
Balance Sheet
Assets Amount($) liabilities Amount($)
Cash 3500 Accounts Payable 600
Accounts Receivable 8000 Income Tax Pyable 8800
Prepaid Insurance 3600 Stockholders Equity 68700
Equipment                                            75000
Less: Accumulated Depreciation 12000 63000
78100 78100

A) Current Ratio = Current Assets / Current liabilities= 3500+8000/600+8800 =11500/9400=  1.2234

B) Working Capital = Current Assets - Current liabilities ==11500-9400 =2100

C) Assets Turnover= Total Sales / average of Total Asses

  average of Total Asses =3500+8000+3600+63000/4 = 19525

Assets Turnover= 102200/19525= 5.234315

D) Fixed Assets Turnover= Net Sales/ Average Fixed Assets= 102200/63000= 1.622222

E). Accounts Receivable Turnover =Net Sales/ Average Accounts Receivables = 102200/8000= 12.775

F). Debt Ratio = Total Liabilities / Total Assets = 9400/78100 = 0.120358